# An Ising Model of Bursts and Fizzles in Intergroup Violence

### v8

@@ 4.1 Operationalization
We coded violence when force was used against another’s body (punching, slapping, kicking, hitting, stomping) and/or when another person’s body was forcefully moved. People may defect (not use violence) for all kinds of reasons and due to various causes. Yet, unconditional defectors have in common that they are (temporarily) inactive while others contribute. Among them we include deescalators, as well as group members who were prevented to fight by deescalators.
@@
We subsumed the following behaviors of members of the opponent group under turmoil for the focal group: aggressing, including fighting gestures; pulling off clothing (jackets or vests); pulling up pants; pointing toward opponents; provocative gesturing with fingers or hands (as an invitation to engage); bending forward toward an opponent; encroaching (invading opponents’ personal space through using or damaging objects belonging to them); teasing, such as lightly hitting or ridiculing; and violence. Because over small time intervals, turmoil accumulates in participants’ minds, we calculated the level of turmoil from the beginning of the video until a focal group’s (first) maximum participation in violence, by multiplying the duration of each instance of turmoil by the number of individuals involved, and adding up all these weighted instances. This scale should be improved in future studies, but our main prediction is independent of it.
@@
Given the distraction caused by turmoil, it is not feasible for all group members to react within 1 second to a first cooperator, as in a well-organized sports team, whereas 3 seconds is too long for the notion of burst to apply to our small groups. Hence we defined a burst as an outbreak of violence by at least half of the group (Fig. 2), or both individuals in a dyad, if they started fighting less than 2 seconds after the first, with a 5% margin.
@@
Figure 3: The proportion of unconditional defectors, $p$, and the predicted critical level $p_{c}$ (based on $C=1$ and $D=1/2$) as vertical line, with box-plots of bursts (1) and fizzles (0).
4.2 Empirical results
Of the 59 groups considered, there were 23 groups where violence started in a burst, 15 groups where violence was collective without a burst, and 21 cases of violence by a single group member, thus no burst either. The proportions of defectors in groups with bursts ($\bar{p}=0.19$; sd = 0.21) and groups without bursts ($\bar{p}=0.49$; sd = 0.27) are box-plotted in Fig. 3 (Welch test $t=4.796;P=6.411\cdot 10^{-6}$; df = 54.76). The critical level of $p_{c}=1/3$, drawn as vertical line in Fig. 3, separates most of the bursts from the nonbursts, although some groups are on the wrong side of the line.
Turmoil preceded all collective violence with one exception, where two individuals suddenly assaulted a passive victim. As said, we have no accurate scale for turmoil, and our assessment is tentative. Moreover, $T_{c}$ is case-specific, co-depends on group size, both in absolute number [5] and relative to the size of the opponent group, and increases with $p$ (up to $p_{c}$). Consistent with simulations [5], we found that cooperation in small groups tends to start at lower turmoil than in larger groups (cor. = 0.53) and in the latter, it tends to start in small subgroups. This bottom up mounting of cooperation is similar to bottom up synchronization in the Kuramoto model [2]. The size-turmoil relation is slightly disturbed by dyad-groups more often facing larger opponent groups, and therefore being less likely to fight than triads. As a matter of fact, larger groups always defeated smaller opponent groups (made them flee or worked them to the ground) with only one exception.
5 Conclusions
With asymmetric spins, the Ising model explains when ($T_{c}$), how (Metropolis), and in what pattern (burst/fizzle) collective violence commences, without recourse to initiative takers [17], perfect information [23], or traditional explanations based on reputations [27], prosocial norms [13], or feedback through selective incentives [28]. The traditional mechanisms are recent in evolutionary history, whereas agitating stimuli are billions of years old. Therefore the model might also be useful in biology, for example to study quorum sensing bacteria [11]. The model can be applied to cooperation in other domains of human collective action, such as street protests [16] or helping victims [31]. For different domains, specific scales of agitating stimuli will have to be developed, which will be easier for lab experiments than field studies. If required, the model can incorporate norms (as external fields), selective incentives (through individualized $S_{0,i}$), and noisy reputations [4]. In all likelihood, more discoveries lay ahead.
Data and code accessibility
All coded videos and the R-script used to produce graphs thereof are available at https://osf.io/f25nq/
@@ 5 Conclusions
Data and compilation of data set
@@
One concern about phone-recorded clips is that recorders start filming when the antagonism is already ongoing; our data are thus left-truncated to an unknown degree. We discarded footage that immediately started with physical violence. Another concern is that uploaded videos might be biased toward more spectacular cases. However, our sample shows variety in the forms and severity of violence, ranging from groups who engage in frenzied collective violence to incidents that involve just a few slaps. A third concern is whether the recording influenced the behavior of the assailants. Several videos showed more than one person recording the incident. Earlier work [jones85] noted that a host of research indicates that people do not change their behavior substantially in the presence of a camera and mainly focus on what they are doing. This is probably even more the case in antagonistic situations, and because young people are used to being filmed by phones. Additionally, when people do pay attention to the camera, which happened in just one of our clips, it is visible in the recording and can be incorporated in the analysis. (In our case, this concerned behavior coded as turmoil.)
@@
Ethics
The use of videos for research purposes poses distinct ethical challenges, largely due to the non-anonymous content of the videos. However, ethical guidelines for digital spaces tend to be less restrictive [25], with the consent of the participants being less stringent for data acquired from the public domain, including the internet. While our video corpus is open to use and inspection by other researchers upon request, we require that they take the same measures to ensure the anonymity of the persons portrayed as we did.
Coding
@@
To identify whether individuals belonged to a focal, opponent, or third-party group, we observed whether they remained close (at touching distance) to each other as they moved through space, whether they called each other by their names, whether they touched each other (for instance, by tapping shoulders), or whether they shared a vehicle or other object (for instance, a bag or skateboard). One of the authors and two assistants discussed each case in which we observed one or more of these behaviors to determine whether the actors were part of the same group.
Composition of indicators
@@
Turmoil. We subsumed the following behaviors of members of the opponent group under the heading of turmoil for the focal group. First, we coded approaching when opponents moved closer to focal group members. We also indicated whether actors moved energetically (jumping/skipping) and whether they ran. Second, aggressive behaviors when actors in the opponents’ group moved body parts in a way that signaled provocation, hostility and/or a readiness to attack. We used the following modifiers (sub-codes) to specify aggressing behavior: fighting gestures; pulling off clothing (jackets or vests); pulling up pants; pointing at opponents; provocative gesturing with fingers or hands (as an invitation to engage); bending forward (head and/or chest bending toward opponent); encroaching (invading opponents’ personal space through using or damaging objects belonging to them); and teasing (invading opponents’ body space in a teasing/ridiculing way that does not inflict bodily harm, such as lightly hitting, stroking, or patting the opponent’s body). Third, involuntary behavior that increases the vulnerability of opponents, because it tends to agitate (signaling opportunity) and provoke violence. This concerned stumbling and falling to the ground. Finally, violence committed by members of the opponent group. Because we do not know the psychological impact of the various kinds of turmoil, we could not construct a ratio level scale. Instead, we calculated the total level of turmoil from the beginning of the clip until a focal group’s maximum participation in violence by noting the duration of each instance of turmoil, and multiplying it by the number of opponents involved, i.e., the surface area under the turmoil curve in Fig. 7, 10, and S11.
@@
Bursts. We identified bursts when, at the first moment of collective violence, two conditions are fulfilled: (1) at least half of the focal group members joined the fight, or both actors did in a dyad, and (2) cooperators started fighting in less than 2 seconds (with a 5% margin) after the first. Condition (1) follows from simulations of small groups; see Table 3 and Fig. 13. Note that in simulations of much larger networks ($n\geq 1000$), bursts can occur wherein less than half of a group participates. In some videos, we observed multiple bursts; in these cases, we only considered the first burst to facilitate comparison of cases.
Alternative explanation of collective action
A plausible alternative explanation of the onset of collective action is synchrony of motion [24], which yields a feeling of oneness among group members [14] and a stronger willingness to take risks for one’s group mates [33].
We measured the synchronization of focal group members pairwise if they engaged in the following behaviors simultaneously (i.e., overlapping intervals on the timeline). Approaching members of the opponent group at the same pace (normal, energetic, or running), distancing from the opponent group at normal walking speed (we consider simultaneous distancing at energetic or running pace uncontrolled attempts to escape rather than forms of synchrony), or simultaneously engaging in any of the forms of aggressing (above). To calculate the level of synchrony, we noted the duration of each instance of synchrony from the beginning of the clip prior to the moment of maximum participation in violence, and we multiplied it by the proportion of focal group ties involved (i.e., normalized with respect to the maximum number of ties in the group).
Although in 21 out of 23 bursts, some degree of synchronization preceded collective violence, there were 18 cases in which synchronization was not followed by collective violence. In some of the latter cases, synchrony turned out to be a deceptive performance composed of blustering and aggrandizing without commitment to fighting. This does not imply that synchronization is unimportant (it probably is for solidarity [9]), but it just does not predict collective violence.
Examples from videos
@@
Graphs of behavior on the timeline enable to analyze and compare the coded videos. Each graph shows the intervals and their durations of (collective) violence, turmoil, synchrony and deescalation (see Fig. 7, 10, 11).
@@
To illustrate, we discuss how coded video data was turned into a graph for clip 26. Fig. 4-6 feature screenshots of the clip, and Fig. 7 displays the graph. This clip features a confrontation of a dyad (focal group) with a triad (opponents) in a restaurant. Violence started at approximately 14 seconds into the video and quickly became collective until approximately 28 seconds. Four instances of turmoil preceded the violence, starting with one of the members of the opponent group gesturing aggressively and walking around in a jumpy, energetic way. The members of the focal dyad walked synchronously toward and away from the opponent in three shorter and one relatively longer stretch of time. Fig. 4 captures a moment of synchrony; the members of the dyad were making fighting gestures and stood aligned, each stretching out one leg with their feet pointing toward the opponent, and the other leg standing backward. Fig. 5 shows the moment when the dyad’s level of agitation reached a critical level; when one of the focal actors dodged backward, avoiding being hit, his companion was about to deliver a blow to the head of the opponent, which made the latter fall down. Fig. 6 displays the second member of the opponents’ triad (soon joined by the third) rushing into the restaurant toward the focal dyad, near his fallen companion. At this point, both groups were engaged in collective violence. Fig. 5 also shows bystanders (who can be seen sitting at a table in Fig. 4) and employees standing near the counter. At 28 seconds, an employee took deescalatory action for the first time. At that point, the collective violence had stopped already and erupted again for a shorter moment approximately 2 seconds later. The clip ends when two triad members carry their fallen group member outside. By then, the focal dyad had already left the scene.
@@
Fig. 8-10 provide another illustration. Fig. 10 is the graph of clip 86; it shows three outbreaks of collective violence in a group of five members who faced a single opponent. In the first instance of collective violence (not a burst by our 2 seconds criterion), starting at approximately 60 seconds into the clip and lasting for approximately 30 seconds, four members participated. Three of them immediately followed each other, and the fourth joined after approximately 2 seconds. Prior to the outbreak of violence, the opponent agitated focal group members by walking back and forth and by posturing aggressively. Fig. 8 captures the trigger moment of agitation; i.e., while one member of the focal group was engaged in aggressive posturing and a group member tried to deescalate, the opponent was about to strike another focal group member. Fig. 9 shows the situation 4 seconds later, when three members of the focal group were attacking.
@@
Analysis of coded and plotted videos
We used graphs of coded videos in our analysis, illustrated in Fig. 11, where the opponent group in clip 67 is plotted. The horizontal axis is the timeline in thousands of seconds, the vertical axis is the proportion of group members involved (for turmoil and violence), and the proportion of intragroup ties involved out of the maximum possible number of ties (for synchrony of action). The first moment of collective violence, denoted by the downward arrow, involves two focal group members. There is, however, a short interruption of collective violent action. Therefore, we did not code this occurrence of collective violence as a burst. The time gap between the first participant and maximum participation in collective violence, denoted by the horizontal arrows, takes over 6 seconds.

### v9

@@ 3 Data and Methods
We coded violence when force was used against another’s body (punching, slapping, kicking, hitting, stomping) and/or when another person’s body was forcefully moved. People may defect (not use violence) for all kinds of reasons and due to various causes. Yet, unconditional defectors have in common that they are (temporarily) inactive while others contribute. Among them we include de-escalators, as well as group members who were prevented to fight by de-escalators.
@@
We subsumed the following behaviors of members of the opponent group under turmoil for the focal group: aggressing (including fighting gestures); pulling off clothing (jackets or vests); pulling up pants; pointing toward opponents; provocative gesturing with fingers or hands (as an invitation to engage); bending forward toward an opponent; approaching the focal group; encroaching (invading opponents’ personal space through using or damaging objects belonging to them); teasing, such as lightly hitting or ridiculing; and violence. Because stumbling and falling because vulnerability of opponents tends to agitate focal group members [15, 40, 62], we included these mishaps in our measure. It is likely that over relatively short time intervals, turmoil accumulates in participants’ minds. Therefore we calculated the level of turmoil from the beginning of the video until a focal group’s (first) maximum participation in violence. We did this by multiplying the duration of each instance of turmoil by the number of individuals involved, and adding up all these weighted instances. This scale should be improved in future studies, but our main prediction is independent of it.
@@
Given the distraction caused by turmoil, it is not feasible for all group members to react within 1 second to a first cooperator (as they might in a well-organized sports team), whereas 3 seconds is too long for the notion of burst to apply to our small groups. Hence we defined a burst as an outbreak of violence by at least half of the group (Fig. 3), or both individuals in a dyad, if they started fighting less than 2 seconds after the first, with a 5% margin. [^9] The requirement that $N_{C}\geq 0.5n$ is based on simulations of small networks just above $p_{c}$, but under this condition in large networks (e.g., $n=1000$), $N_{C}<0.5n$.
4 Results
Of the 59 groups considered, there were 23 groups where violence started in a burst, 15 groups where violence was collective without a burst, and 21 cases of violence by a single group member. Turmoil preceded all collective violence with one exception, where two individuals suddenly assaulted a passive victim. The critical level of turmoil ($T_{c}$) for bursts is case-specific and depends on group size, both in absolute number and relative to the size of the opponent group, and on the proportion of unconditional defectors. Additionally, the use of weapons has an intimidating effect.
We confirm the finding of an earlier study [15] that fighting tends to start in small groups or in small subgroups of larger groups. As predicted, we found that small (sub)groups burst into action at lower levels of turmoil than larger groups [9], and lower levels are of course reached earlier. In our data, bursts developed in 13 (37%) of the 35 smallest groups (dyads and triads) and in 10 (42%) of the 24 larger groups. In bursts, the correlation between group size and turmoil is $0.53$. The size-turmoil relationship is slightly disturbed by dyads more often facing a larger opponent group and being therefore less likely to fight collectively than triads, which in turn are more robust in general [52]. In fact, larger groups always defeated smaller opponent groups (made them flee or worked them to the ground) with only one exception.
The proportions of unconditional defectors in groups with bursts (mean = 0.19; s.d. = 0.21) and groups without bursts (mean = 0.49; s.d. = 0.27) are box-plotted in Fig. 5 (Welch test $t=4.796;P=6.411$ x $10^{-6}$; df = 54.76). The predicted critical threshold (vertical line in the figure) neatly separates the two boxplots, but one might question how robust this result is, given some noise in coding. There is a risk of missing a violent act (it can be very fast) or mislabeling a fast movement as violence, thus mixing up fighters and defectors, and a risk of misassigning individuals to groups. Even if we suppose there is a 10% chance of each of these errors (which corresponds to Krippendorff’s $\alpha=0.75$; see inter-coder reliability in the Supplementary Material), the result is hardly weakened (Welch test $t=3.879;P=0.0025$, averaged over 1000 simulation runs), hence the distinction between bursts and fizzles appears to be robust.
@@ 4 Results
Figure 5: The proportion of unconditional defectors and the predicted critical level (vertical line), which for our data separates the boxplots of bursts (1) and fizzles (0).
The empirical distinction between bursts and fizzles is not perfect, and Fig. 5 shows that 15 groups out of 59 (25%) are categorized incorrectly. In 11 of these 15 groups, the time between the first violent act and last participant joining in was beyond our two-second interval, hence we did not count these collective actions as bursts. In simulations, lags between first instances of violence and the moment when at least half of the group is participating frequently occur when turmoil lingers at its critical level (Fig. 6), versus steadily increasing turmoil that yields a clean burst (Fig. 3). Lingering turmoil seems to have occurred in 4 cases, but we would need a better scale of turmoil than the cumulative scale we currently use to be sure. Far more important are local circumstances beyond the model. When re-examining the time lags beyond 2 seconds, we noticed various causes of delay: (1) group members were in a car and it took time to get out; (2) a focal group member stood at a distance from the fight on a slippery floor next to a pool; (3) a focal group member was beaten and seemed intoxicated by alcohol, hence responded slowly; (4) a focal group member was first pushed over and it took him time to get on his feet again; (5) opponents moved around so fast that it took time to hit one; (6) an opponent fell into thick bushes such that only one focal group member at a time could get to him; (7) group members were sitting in a train and needed time to get on their feet; (8) group members were constrained in a small room where they searched for a space to lash out; and, (9) group members were hindered by the mess of two opponents fighting on a collapsed table full of food. One can hardly blame the model for not incorporating these local circumstances that caused delays, and perhaps our 2 seconds limit should be applied with more care. Finally, two delays were caused by ambiguous behavior: (10) someone tried to de-escalate, was hit, and then attacked, and (11) someone tried to pull back his fighting friend (de-escalation) but attacked simultaneously.
Figure 6: Proportion of cooperators ($N_{C}/n$) over time (or precisely, over Monte Carlo steps) in a fully connected network ($n=5$) with one unconditional defector, at $T=0.11\approx T_{c}$. Note an extended period of multiple short spikes of violence before all conditional cooperators join in. Similar lags occur without unconditional defectors.
In the 4 out of 15 remaining groups, non-bursts were predicted whereas bursts occurred. Although for parsimony, we applied the same $S_{0}$ (and $\Delta$) for all conditional cooperators, some combatants ($i$) may actually have a higher value than the average, $S_{0,i}>\overline{S_{0}}$, which could explain these cases. In lab experiments, it is unlikely that the patterns of bursts and fizzles would have been discovered, but now they have been pointed out in the field, ethically acceptable intergroup conflicts in the lab may benefit from controlled circumstances, such as equal monetary payoffs for all cooperators, and may lead to refinements of the model.
4.1 Alternative explanation
A plausible alternative explanation for the onset of collective action is synchrony of motion [37], which yields a feeling of oneness among group members [20] and a stronger willingness to take risks for one’s group mates [56].
We measured the synchronization of focal group members pairwise if they engaged in the following behaviors simultaneously (i.e., overlapping intervals on the timeline): approaching members of the opponent group at the same pace (normal, energetic, or running), distancing from the opponent group at normal walking speed (we consider simultaneous distancing at energetic or running pace uncontrolled attempts to escape rather than forms of synchrony), or simultaneously engaging in aggressing. To calculate the level of synchrony, we noted the duration of each instance of synchrony from the beginning of the clip prior to the moment of maximum participation in violence, and we multiplied it by the proportion of focal group ties involved (i.e., normalized with respect to the maximum number of ties in the group).
Although in 21 out of 23 bursts, some degree of synchronization (10.8 on average) preceded collective violence, there were 18 cases in which synchronization (9.6 on average) was not followed by collective violence. In some of the latter cases, synchrony turned out to be a deceptive performance composed of blustering and aggrandizing without commitment to fighting. This does not imply that synchronization is unimportant (it probably is for solidarity [15]), but it does not predict collective violence.
5 Discussion
The simple Ising model is a century old [31] and has been applied to a wide range of problems [44, 55], to which we add the dilemma of collective action. It explains cooperation parsimoniously, based on agitating stimuli without recourse to strong rationality, initiative takers, reputations, norms, feedback through selective incentives, or reliable information passing through the network. The Ising model elucidates the temporal unfolding of violence by predicting a critical threshold of unconditional defectors that distinguishes a burst of collective action from a fizzle, which is largely supported by the data. The model also explains why violent groups are often small or are small subgroups of larger groups despite greater risk. Small (sub)groups have a lower critical threshold of turmoil, and in a confrontation with opponents, lower levels are reached earlier. We also investigated whether synchronous action precedes violence, but we found that synchronization precedes both collective and solitary violence, and cannot predict either of these outcomes. However, synchronization may still be important to increase solidarity [18].
This study has several limitations. Because we did not select any videos without violence, we cannot be certain that violence is caused by turmoil. When investigating the videos, however, we observed time and again that people reacted violently to turmoil, hence it still seems quite plausible that turmoil triggers violence. Moreover, the model predicts that when turmoil increases, small groups start cooperating earlier than large groups, which we were unable to falsify. Our measurements underestimated turmoil because the videos are left-truncated, and our data depend on camera angle and vision width, inevitably excluding instances of turmoil. Furthermore, there might be some error in our coding. If the truth were perfectly known and the coding corrected accordingly, the two boxplots in Fig. 5 might shift to a small degree, but this would would not affect our main result. In the future, a better scale of agitation should be developed. The scale should differentiate between domains (e.g., street violence versus grievances about politics), thus there should be multiple scales, and these should differentiate between effect sizes of different stimuli. For example, the impact of family members being killed will be much larger and last longer than that of a push, which might be quickly forgotten. Some protest movements are known to produce agitation by themselves, which is sometimes indicated by accelerated posting of online messages. Because in these cases, the inter-posting time intervals follow a certain pattern (Moore’s law), $T_{c}$ can be accurately predicted, well ahead of large street demonstrations [28]. Another limitation is in the Ising model itself. Despite predicting the threshold of unconditional defectors well, some of the groups ended up in the non-burst category due to local circumstances (ignored by the model) that caused delays beyond our two second limit. Furthermore, the Ising model does not predict the severity of violence. For future studies, it is important to expand the number and diversity of cases, and to develop better software than is currently available to automatically code videos. Finally, when individuals find themselves more often in similar situations, they will learn, which is easier in smaller groups where they have a larger influence on their payoff [11]. Some will change their strategy, and turn into unconditional defectors [2] who try to exploit other group members and maximize their individual payoff instead of maximizing the group’s payoff. For the model, this would require individuals updating their decision rules at subsequent Monte Carlo steps.
Here we applied the Ising model to violence, but it can also be applied to other uncertain situations. Its dynamics are entirely consistent with temporal patterns of protests [24], which break out more often if (rumors say that) a government or its police are weakened [53, 58], analogous to vulnerable individuals in street fights. The model also seems applicable to spontaneous lynchings [6], and bystanders collectively helping victims under uncertainty [47]. In this first empirical application, we showed that it can explain the unfolding of street violence, and in all likelihood, more discoveries lay ahead. Extensions to norms (as external fields) and noise in actors’ information about others’ behavior have been explored in simulations [8].
Taking a broader perspective, random noise can solve coordination problems, such as collective responses to opponents’ threats, by shaking a group loose from its suboptimal (e.g., non-cooperative) state [51], but it can also disturb an optimal state. In the Ising model, both can happen, depending on the amount of turmoil: when in a group, turmoil approaches $T_{c}$ from below, random noise in decision making facilitates cooperation, but at high levels, too much noise entails confusion.
Ethics
The use of videos for research purposes poses distinct ethical challenges, largely due to the nonanonymous content of the videos. However, ethical guidelines for digital spaces tend to be less restrictive [38], with the consent of the participants being less stringent for data acquired from the public domain, including the internet. While our video corpus is open for use and inspection by other researchers upon request, we require that they take the same measures to ensure the anonymity of the persons portrayed as we did.
Data availability statement
All coded video data, the R script used to produce plots (Supplementary Material) from the data, and a Fortran script for simulations of the Ising model are available at https://osf.io/f25nq/
@@ 5 Discussion
6 Data and compilation of data set
@@
One concern about phone-recorded clips is that recorders start filming when the antagonism is already ongoing; our data are thus left-truncated to an unknown degree. We discarded footage that immediately started with physical violence. Another concern is that uploaded videos might be biased toward more spectacular cases. However, our sample shows variety in the forms and severity of violence, ranging from groups who engage in frenzied collective violence to incidents that involve just a few slaps. A third concern is whether the recording influenced the behavior of the assailants. Several videos showed more than one person recording the incident. Earlier work [30] noted that a host of research indicates that people do not change their behavior substantially in the presence of a camera and mainly focus on what they are doing. This is probably even more the case in antagonistic situations, and because young people are used to being filmed by phones. Additionally, when people do pay attention to the camera, which happened in just one of our clips, it is visible in the recording and can be incorporated in the analysis. (In our case, this concerned behavior coded as turmoil.)
@@
7 Coding
@@
[…]
8 Composition of indicators
@@
Turmoil. We subsumed the following behaviors of members of the opponent group under the heading of turmoil for the focal group. First, we coded approaching when opponents moved closer to focal group members. We also indicated whether actors moved energetically (jumping/skipping) and whether they ran. Second, aggressive behaviors when actors in the opponents’ group moved body parts in a way that signaled provocation, hostility and/or a readiness to attack. We used the following modifiers (sub-codes) to specify aggressing behavior: fighting gestures; pulling off clothing (jackets or vests); pulling up pants; pointing at opponents; provocative gesturing with fingers or hands (as an invitation to engage); bending forward (head and/or chest bending toward opponent); encroaching (invading opponents’ personal space through using or damaging objects belonging to them); and teasing (invading opponents’ body space in a teasing/ridiculing way that does not inflict bodily harm, such as lightly hitting, stroking, or patting the opponent’s body). Third, involuntary behavior that increases the vulnerability of opponents, because it tends to agitate (signaling opportunity) and provoke violence. This concerned stumbling and falling to the ground. Finally, violence committed by members of the opponent group. Because we do not know the psychological impact of the various kinds of turmoil, we could not construct a ratio level scale. Instead, we calculated the total level of turmoil from the beginning of the clip until a focal group’s maximum participation in violence by noting the duration of each instance of turmoil, and multiplying it by the number of opponents involved, i.e., the surface area under the turmoil curve in Fig. S10, S13, and S14.
@@
Bursts. We identified bursts when, at the first moment of collective violence, two conditions are fulfilled: (1) at least half of the focal group members joined the fight, or both actors did in a dyad, and (2) cooperators started fighting in less than 2 seconds (with a 5% margin) after the first. Condition (1) follows from simulations of small groups; see Table S3. Note that in simulations of much larger networks ($n\geq 1000$), bursts can occur wherein less than half of a group participates. In some videos, we observed multiple bursts; in these cases, we only considered the first burst to facilitate comparison of cases.
9 Examples from videos
@@
Graphs of behavior on the timeline enable to analyze and compare the coded videos. Each graph shows the intervals and their durations of (collective) violence, turmoil, synchrony and deescalation (see Fig. S10, S13, S14).
@@
To illustrate, we discuss how coded video data was turned into a graph for clip 26. Fig. S7-9 feature screenshots of the clip, and Fig. S10 displays the graph. This clip features a confrontation of a dyad (focal group) with a triad (opponents) in a restaurant. Violence started at approximately 14 seconds into the video and quickly became collective until approximately 28 seconds. Four instances of turmoil preceded the violence, starting with one of the members of the opponent group gesturing aggressively and walking around in a jumpy, energetic way. The members of the focal dyad walked synchronously toward and away from the opponent in three shorter and one relatively longer stretch of time. Fig. S7 captures a moment of synchrony; the members of the dyad were making fighting gestures and stood aligned, each stretching out one leg with their feet pointing toward the opponent, and the other leg standing backward. Fig. S8 shows the moment when the dyad’s level of agitation reached a critical level; when one of the focal actors dodged backward, avoiding being hit, his companion was about to deliver a blow to the head of the opponent, which made the latter fall down. Fig. S9 displays the second member of the opponents’ triad (soon joined by the third) rushing into the restaurant toward the focal dyad, near his fallen companion. At this point, both groups were engaged in collective violence. Fig. S8 also shows bystanders (who can be seen sitting at a table in Fig. S7) and employees standing near the counter. At 28 seconds, an employee took deescalatory action for the first time. At that point, the collective violence had stopped already and erupted again for a shorter moment approximately 2 seconds later. The clip ends when two triad members carry their fallen group member outside. By then, the focal dyad had already left the scene.
@@
Fig. S11-S13 provide another illustration. Fig. S13 is the graph of clip 86; it shows three outbreaks of collective violence in a group of five members who faced a single opponent. In the first instance of collective violence (not a burst by our 2 seconds criterion), starting at approximately 60 seconds into the clip and lasting for approximately 30 seconds, four members participated. Three of them immediately followed each other, and the fourth joined after approximately 2 seconds. Prior to the outbreak of violence, the opponent agitated focal group members by walking back and forth and by posturing aggressively. Fig. S11 captures the trigger moment of agitation; i.e., while one member of the focal group was engaged in aggressive posturing and a group member tried to deescalate, the opponent was about to strike another focal group member. Fig. S12 shows the situation 4 seconds later, when three members of the focal group were attacking.
@@
10 Analysis of coded and plotted videos
We used graphs of coded videos in our analysis, illustrated in Fig. S14, where the opponent group in clip 67 is plotted. The horizontal axis is the timeline in thousands of seconds, the vertical axis is the proportion of group members involved (for turmoil and violence), and the proportion of intragroup ties involved out of the maximum possible number of ties (for synchrony of action). The first moment of collective violence, denoted by the downward arrow, involves two focal group members. There is, however, a short interruption of collective violent action. Therefore, we did not code this occurrence of collective violence as a burst. The time gap between the first participant and maximum participation in collective violence, denoted by the horizontal arrows, takes over 6 seconds.